1) If ​$4000 is deposited in a savings account that earns interest at an annual rate of 2.5​% interest compounded​ continuously, what is the value of the account at the end of two ​years?

2) A trust fund for a 11​-year-old child is being set up by a single payment so that at age 21 the child will receive ​$37,000. Find how much the payment is if an interest rate of 9​% compounded semiannually is assumed.

3) A bank account pays 5.5​% annual​ interest, compounded monthly. How much must be deposited now so that the account contains exactly ​$12,000 at the end of one year​?
Answer #1
1) Value of account at the end of 2 years = P * e^(i*n) Where,
= $   4,000.00 * 2.7183^(0.025*2) P = $         4,000.00
= $   4,000.00 * 1.051271 e = 2.7183 (A mathematical constant)
= $   4,205.09 i = 2.50%
n = 2
2) Present value = P * (1+i)^-n Where,
= $ 37,000.00 * (1+0.045)^-20 P = $      37,000.00
= $ 37,000.00 * 0.414643 i = 4.50%
= $ 15,341.79 n = 20
3) Present value = P * (1+i)^-n Where,
= $ 12,000.00 * (1+0.0045833)^-12 P = $      12,000.00
= $ 12,000.00 * 0.946604 i =          0.0045833
= $ 11,359.25 n = 12